# spacing=nonfrench

### v1

@@ 7. Related Work
AutoTuning. AutoTuning allows developers to automatically find the best values of various configuration parameters at runtime without requiring the developers to manually try all possible combinations. Similar to FDO techniques, AutoTuning techniques (Van Aken et al., 2017; Sriraman and Wenisch, 2018) often perform tuning through trial runs to find optimal values for different parameters. Some AutoTuning techniques such as OPPerTune (Somashekar et al., 2024) and SoftSKU (Sriraman et al., 2019) apply the AutoTuning paradigm of measurement-driven search at runtime without any trials. However, these techniques are limited to environmental parameters and knobs as they work transparently with respect to the deployed application. Iridescent enables the AutoTuning paradigm of measurement-driven search at runtime to find the best combination of specializations for the system.
8. Discussion
@@
Debugging. Iridescent makes it harder for developers to issues. This is because the generated specialized code is often very different to the code that was originally written by the developer which makes it harder for the developers to fully understand the execution sequences. This is further exacerbated by the fact that different workloads may lead to different specializations making it difficult to reproduce issues seen in production. Combining Iridescent with monitoring and distributed tracing techniques could offer developers a solution to their debugging problems.
Profile-Guided Optimizations at runtime. Iridescent enables developers to utilize traditional PGO techniques at runtime without requiring separate benchmarking runs. For example, prefetch injection optimizations such as APT-GET (Jamilan et al., 2022) and RPG2 (Zhang et al., 2024) could be integrated with Iridescent in the future.
Leveraging scale & ML for exploration. Several ML-based techniques exist for tuning configurations (Van Aken et al., 2017; Akgun et al., 2023; Cereda et al., 2021; Li et al., 2018; Lin et al., 2022; Mahgoub et al., 2020; Somashekar et al., 2022, 2024; Duan et al., 2009; Jamilan et al., 2022; Swaminathan et al., 2017; Ruffy et al., 2024). OPPerTune (Somashekar et al., 2024) provides access to a plethora of ML techniques toolbox for automatically searching through large configuration spaces. We believe Iridescent can be integrated with OPPerTune to utilize state-of-the-art exploration techniques. Moreover, Iridescent could further incorporate SmartChoices (Carbune et al., 2018) to allow different ML techniques to apply to specific specialization points.
9. Conclusions
@@
In this work, we presented Iridescent, a framework enabling online workload-driven runtime specialization of systems to improve performance. Iridescent provides a toolkit for developers to implement system-specific specializations that utilize runtime data and invariants to generate more performant systems. Iridescent provides the necessary tooling for developers to configure automatic runtime exploration of the space of potential specializations to find the best performing configuration at each instant. By combining developer insight with automated compiler optimizations, Iridescent enables systems to adapt efficiently to changing workloads and environments with minimal manual effort to improve performance.

### v2

@@ 7.2 Online Automatic Optimizations
Runtime auto-tuning techniques such as OtterTune [82] often perform tuning through trial sessions to find optimal values for different configuration knobs. Some AutoTuning techniques such as OPPerTune [75] and SoftSKU [77] apply the AutoTuning paradigm of measurement-driven search at runtime without any trials. However, these techniques are critically limited to environmental parameters and knobs as they work transparently with respect to the deployed application. These techniques are currently not applicable to code optimizations or specializations. Iridescent enables the auto-tuning paradigm of measurement-driven search at runtime to find the best combination of code specializations for the system.
Code specializations.
Runtime code specialization techniques [18, 17, 32, 22, 5, 65] produce more efficient versions of the code by exploiting values and invariants that only exist at runtime. Today, these specializations manifest in three ways: (i) binary rewriting techniques [3]; (ii) Transparent Dynamic Optimization (TDO) techniques such as Dynamo [6], The Transmeta Code Morphing Software [23], DynamoRIO [9]. These optimize code at runtime without requiring any modifications by capturing and optimizing traces (sequences of instructions) that are commonly executed; (iii) optimizations in JIT compilers for interpreted languages in the form of Type Specialization [15, 57, 14] for specializing the types of data for dynamically typed languages or as Value Specialization for interpreted languages [21, 49] where the parameter values of hot functions are converted into constants. These specializations may be optionally applied depending on the computation context [37, 63, 28, 54]. In contrast to Iridescent, they are generic and typically done automatically by the runtime based on internal cost models which may not be appropriate for specific application contexts. Moreover, these specializations are limited in scope and do not take the system’s end-to-end performance into consideration and do not leverage domain- and situation-specific optimizations.
Domain-specific specializations.
Recently, value specializations have extended beyond interpreted languages for specific use cases such as value specializations for GPU kernels [29], or through incremental specializations for packet processing frameworks [68, 53, 27]. Iridescent provides a general framework for configuring, applying, exploring, and selecting such specializations at runtime.
8 Discussion
Debugging.
@@ 8 Discussion
Iridescent makes it harder for developers to debug performance issues. This is because the generated specialized code is often different from the code that was originally written by the developer, which makes it harder for the developers to fully understand the execution sequences. This is further exacerbated by the fact that different workloads may lead to different specializations making it difficult to reproduce issues seen in production. Combining Iridescent with monitoring and distributed tracing techniques could offer developers a solution to their debugging problems.
Leveraging scale & ML for exploration.
Several ML-based techniques exist for tuning configurations [82, 1, 12, 48, 50, 51, 74, 75, 26, 38, 80, 68]. These ML techniques can be implemented as algorithms with OPPerTune [75], which we have already integrated with Iridescent. Moreover, Iridescent could further incorporate SmartChoices [10] to allow different ML techniques to apply to specific specialization points. As part of future work, Iridescent could potentially also parallelize exploration across a large number of replicas to explore larger subsets of the specialization space.
Limitations.
While Iridescent is designed to operate in dynamic and irregular operating conditions, exploration overhead overshadows performance benefits if the operating conditions do not remain stable for long enough to reap the benefits of the specialization optimizations. Ideally, developers should configure the specialization policy to opt for the general version of the handler code in such scenarios.
9 Conclusions
@@
In this work, we presented Iridescent, a framework enabling online workload-driven runtime specialization of systems to improve performance. Iridescent provides a toolkit for developers to implement system-specific specializations that utilize runtime data and invariants to generate more performant systems. Iridescent provides the necessary tooling for developers to configure automatic runtime exploration of the space of potential specializations to find the best performing configuration at each instant. By combining developer insight with automated compiler optimizations, Iridescent enables systems to adapt efficiently to changing workloads and environments with minimal manual effort.